if (H,*) is a subgroup of (G,*) and if R is the relation defined on G by aRb iff a * b^-1 is an element of H, then R is an equivalence relation on G, and that aRb iff b is an element of Ha (where Ha={h * a: h is an element of H}), and thus [a] subscript of R=Ha.Let G=S subscript of 4 , and let H=<{(13),(14)}>. Find H, and find all of the equivalence classes of S subscript of 4 under R.
1
Expert's answer
2020-10-25T18:34:46-0400

Let "(H,*)" be a subgroup of "(G,*)" and if "R" be the relation defined on "G" by "aRb" iff "ab^{-1}\in H".


Show that "R" is an equivalence relation on "G".

Since "a*a^{-1}=e\in H" for any "a\in G", "R" is reflexive relation.

If "aRb", then "a*b^{-1}\in H". Since "H" is a subgroup, "b*a^{-1}=(a*b^{-1})^{-1}\in H". Therefore, "bRa" and "R" is symmetric relation.

If "aRb" and "bRc" , then "a*b^{-1}\in H" and "b*c^{-1}\in H". Since "H" is a subgroup, "a*c^{-1}=(a*b^{-1})*(b*c^{-1})\in H" . Therefore, "aRc" and "R" is transitive relation.

Consequently, "R" is an equivalence relation.


Taking into account that "b*a^{-1}\in H" if and only if there exist "h\in H" such that "b*a^{-1}=h", i.e. "b=h*a\in Ha", we conclude that "bRa" iff "b" is an element of "H*a=\{h*a|h\in H\}". Therefore, "[a]_R=\{b\in G|bRa\}=H*a=\{h*a|h\in H\}".


Consider subgroup "H=\langle{(13),(14)}\rangle=\{(1),(13),(14),(34),(134),(143)\}" of the symmetric group "S_4". Since "|S_4|=24" , there exist "\frac{24}{6}=4" different equivalence classes "S_4" under "R". Let us find them:


"[(1)]_R=H(1)=H=\{(1),(13),(14),(34),(134),(143)\},"


"[(12)]_R=H(12)=\{(1),(13),(14),(34),(134),(143)\}(12)=\{(12), (123), (124), (12)(34), (1234), (1243)\},"


"[(23)]_R=H(23)=\{(1),(13),(14),(34),(134),(143)\}(23)=\{(23), (132), (14)(23), (243), (1324), (1432)\},"


"[(24)]_R=H(24)=\{(1),(13),(14),(34),(134),(143)\}(24)=\{(24), (13)(24), (142), (234), (1342), (1423)\}."